OpenStudy (ohohaye):

How do I find the period for d=9cos(pi/2 t)? Please explain

jimthompson5910 (jim_thompson5910):

\[\Large d = 9\cos\left(\frac{\pi}{2}t\right)\] is the same as \[\Large d = 9\cos\left(\frac{\pi}{2}t-0\right)+0\] Agreed? Or no?

jimthompson5910 (jim_thompson5910):

let me know where you're stuck @ohohaye

OpenStudy (ohohaye):

I'm confused. Can you please explain.

OpenStudy (ohohaye):

I know that the answer is 4 but, I don't know how to get that answer.

jimthompson5910 (jim_thompson5910):

I used the idea that x+0 = x x-0 = x x can be any real number

jimthompson5910 (jim_thompson5910):

the reason I'm doing it is to get the equation into a certain form

OpenStudy (ohohaye):

Are you trying to make the equation to equal 0?

jimthompson5910 (jim_thompson5910):

no, I'm trying to make it fit the form \[\Large y = A*\cos(Bt - C) + D\]

jimthompson5910 (jim_thompson5910):

notice how \[\Large d = 9\cos\left(\frac{\pi}{2}t-0\right)+0\] is in the form \[\Large y = A*\cos(Bt - C) + D\] where A = 9 B = pi/2 C = 0 D = 0 do you see how I'm getting this?

OpenStudy (ohohaye):

Yeah, I see that now

OpenStudy (ohohaye):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

To find the period T, we now use the formula T = 2pi/B

jimthompson5910 (jim_thompson5910):

Plug in B = pi/2 T = 2pi/B T = 2pi/(pi/2) T = 2pi*(2/pi) T = 2*2 T = 4 so the period is 4.